# Approximating the set of Nash equilibria for convex games

### v2

@@ 1 Introduction
In [5], it was shown that the set of Nash equilibria can be equivalently characterized by the set of Pareto optimal points of a specific vector optimization problem with a non-convex ordering cone, or equivalently, by the intersection of the Pareto optimal points of $N$ specific multi-objective problems. This result holds true in general for any possible non-cooperative game without making assumptions on cost functions or constraint sets. In order to use this result for numerical computations, one would need to compute the set of Pareto optimal points for multi-objective problems. So far, algorithms that provide the set of Pareto optimal points exist only for the special case of linear multi-objective optimization, see e.g. [1, 22, 21]. This restricts computational methods of finding the set of Nash equilibria via that characterization to linear games so far.
@@
This result provides some advantages: It is guaranteed that each element of interest, namely each equilibrium point, is contained in the computed set. Furthermore, each element in the found set is guaranteed to be at least almost at equilibrium, i.e. deviating at most by $\epsilon$ from the cost provided by a true equilibrium point. Additionally to the Lipschitz continuity for the convex cost functions, we distinguish between two different assumptions on the structure of the shared constraint set: In the first case we assume that the shared constraint set is a polytope. In the second case we consider for each player $i$ a convex constraint set that is independent from the constraints of the other players. For both cases, the proposed algorithm is proven to work correctly. Computational examples are provided which illustrate the sandwich result.
@@ 2 Definitions
Let us consider the following non-cooperative shared constraint games for $N\geq 2$ players. Each player $i$ (for $i=1,...,N$) considers strategies in the linear space $\mathcal{X}_{i}=\mathbb{R}^{n_{i}}$, where $n_{i}\in\mathbb{N}$, and where we do not impose any condition that $\mathcal{X}_{i}$ and $\mathcal{X}_{j}$ are equal. Furthermore, each player $i$ has a cost function $f_{i}:\prod_{j=1}^{N}\mathcal{X}_{j}\to\mathbb{R}$ she seeks to minimize. This cost $f_{i}(x)$ for $x\in\prod_{j=1}^{N}\mathcal{X}_{j}$ may depend on player $i$’s strategy $x_{i}\in\mathcal{X}_{i}$ as well as the strategy chosen by all other players $x_{-i}\in\prod_{j\neq i}\mathcal{X}_{j}$. The vector of cost functions is denoted by $f=(f_{1},...,f_{N})$ with $f:\prod_{i=1}^{N}\mathcal{X}_{i}\to\mathbb{R}^{N}$. Assume further a shared constraint set $\mathbb{X}\subseteq\prod_{j=1}^{N}\mathcal{X}_{j}$ for the joint strategy $x\in\prod_{j=1}^{N}\mathcal{X}_{j}$ of all players. This shared constraint condition is common in the literature (see, e.g., [18, 3]). In a non-cooperative shared constraint game each player $i$ minimizes her cost function given all other players fix their strategies $x_{-i}^{*}\in\prod_{j\neq i}\mathcal{X}_{j}$. That is, for all $i=1,...,N$ the following optimization problem is considered
@@
When numerical methods are used, one has to deal with approximation errors. Therefore, we will now introduce the notion of approximate Nash equilibria. They are also needed for, e.g., sensitivity analysis of Nash equilibria, see [4].
@@
[17, Section 2.6.6] The joint strategy $x^{*}\in\mathbb{X}$ is called an $\bm{\epsilon}$-approximate Nash equilibrium for the game if, for any player $i$,
$f_{i}(x_{i},x_{-i}^{*})\leq f_{i}(x^{*})\;\Rightarrow\;f_{i}(x_{i},x_{-i}^{*})+\epsilon\geq f_{i}(x^{*})$
for any $x_{i}\in\mathcal{X}_{i}$ such that $(x_{i},x_{-i}^{*})\in\mathbb{X}$. The set of all $\epsilon$-approximate Nash equilibria will be denoted by $\epsilon\operatorname{NE}(f,\mathbb{X})$. That is,
@@
Let us now introduce the notion of Pareto optimality. Consider a linear space $\mathcal{X}$ as well as the space $\mathbb{R}^{m}$ with the natural ordering cone +$\mathbb{R}^{m}_{+}$. Recall that a set $C\subseteq\mathbb{R}^{m}$ is called a cone if $\alpha C\subseteq C$ for all $\alpha\geq 0$. The convex cone +$\mathbb{R}^{m}_{+}$ introduces a pre-order $\leq$ on $\mathbb{R}^{m}$ via $x\leq y\iff y-x\in\mathbb{R}^{m}_{+},$ for $x,y\in\mathbb{R}^{m}$. A multi-objective optimization problem is an optimization problem of the form
@@
We use the notation $g[\mathbb{X}]:=\{g(x)\;|\;x\in\mathbb{X}\}\subseteq\mathbb{R}^{m}$ for the image of the feasible set $\mathbb{X}$. Note that the set $\operatorname*{argMin}\{g(x)\;|\;x\in\mathbb{X}\}$ can equivalently be characterized as the set of feasible points $x^{*}$ that map to minimal elements of $g[\mathbb{X}]$, i.e., the set of points $x^{*}\in\mathbb{X}$ such that if $g(x)\leq g(x^{*})$ for some $x\in\mathbb{X}$ then $g(x^{*})\leq g(x)$ holds, see [7, Def. 3.1(c), Def. 4.1(a) and Def. 7.1(a)].
Denote by $\mathcal{P}:=\operatorname{cl}(g[\mathbb{X}]+\mathbb{R}^{m}_{+})$ the upper image of problem (2). As neither $\mathcal{P}$, nor the set of all Pareto optimal points of problem (2) can be computed exactly in general, except for linear multi-objective optimization problems, one usually computes a (polyhedral) inner approximation of $\mathcal{P}$ for a given error level $\epsilon>0$ and a fixed direction $c\in\mathbb{R}^{m}_{+}\setminus{\{0\}}$, see e.g. [11] for the case of convex problems. This motivated the following definition of $\epsilon$-Pareto optimal points.
@@
[10][6, Def. 2.2] Let $c\in\mathbb{R}^{m}_{+}\setminus{\{0\}}$. An element $x^{*}\in\mathbb{X}$ is called $\bm{\epsilon}$-Pareto optimal with respect to $c$ for problem (2) if
@@
The set of all $\epsilon$-Pareto optimal points of problem (2) with respect to $c$ is denoted by
@@
[6] An element $x^{*}\in\mathbb{X}$ is called weakly Pareto optimal for problem (2) if
$\big{(}g(x^{*})-\operatorname{int}\mathbb{R}^{m}_{+}\big{)}\cap g[\mathbb{X}]=\emptyset.$
Let $c\in\mathbb{R}^{m}_{+}\setminus{\{0\}}$. An element $x^{*}\in\mathbb{X}$ is called weakly $\epsilon$-Pareto optimal with respect to $c$ for problem (2) if
$\big{(}g(x^{*})-\epsilon c-\operatorname{int}\mathbb{R}^{m}_{+}\big{)}\cap g[\mathbb{X}]=\emptyset.$
@@ 3 Equivalence of Nash and Pareto and their approximations
Note that the ordering cone of problem (3) is +$\mathbb{R}^{m_{i}}_{+}$ with $m_{i}=2\sum_{j=1,\\ j\neq i}^{N}n_{j}+1$. This objective space dimension can be further reduced by considering the equivalent convex multi-objective optimization problem
@@
with natural ordering cone in dimension $\sum_{j=1,\\ j\neq i}^{N}n_{j}+2$. This is important for computational aspects and should be kept in mind. For purely aesthetical reasons and since it simplifies explanations we will work with problem (3) instead of problem (4) in our exposition.
@@
The above theorem was given in [5] in a more general framework and in a different formulation using a non-convex ordering cone (see [5, Th. 2.6, Cor. 2.8]). We provide here the above intuitive and simplified formulation that was already mentioned in Remark 3.5 in [5] for the linear case, but holds of course also in general.
@@
Theorems 3.1 and 3.2 can also be stated as:
  • For generalized games, where each player can have an individual constraint set, i.e., when problem (1) in Definition 2.1 is replaced by $x_{i}^{*}\in\operatorname*{arg\,min}\{f_{i}(x_{i},x_{-i}^{*})\;|\;(x_{i},x_{-i}^{*})\in\mathbb{X}_{i}\}$. Then, the constraints sets $\mathbb{X}$ in the Pareto problems in Theorems 3.1 and 3.2 have to be replaced by the individual constraint sets $\mathbb{X}_{i}\subseteq\prod_{j=1}^{N}\mathcal{X}_{j}$ for $i\in\{1,...,N\}$.
  • For general linear spaces $\mathcal{X}_{i}$ for the strategies. Then the ordering cone +$\mathbb{R}^{m_{i}}_{+}$ in problem (3) is replaced by the natural ordering cone in the corresponding space.
@@ 4 Convex games
The aim of this paper is to develop an algorithm that uses representations (5) and (6) to approximate the set of Nash equilibria of a non-cooperative game. To do so, we will focus for the remaining part of this paper on convex games.
@@
As mentioned above in issue (i), we try to cover the set of all Pareto optimal points and therefore make additional assumptions on the structure of the constraint set $\mathbb{X}$. In the following, we will consider constraints sets $\mathbb{X}$ that are polytopes, whereas in Remark 4.11, we consider the case where each player $i$ has an independent convex constraint set $\mathbb{X}_{i}\subseteq\mathcal{X}_{i}$. Problem (iii) can be handled by a small modification of the algorithm in [11] that is possible because of the particular structure of the objective function of problem (3) being linear in all but the last component and the directions $\bar{c}_{i}=(0,...,0,1)^{\top}$ being zero in these linear components.
Let us now discuss problem (ii) in detail. The reason why approximations are considered for convex problems is that it is in general not possible to compute the upper image, but only a polyhedral approximation of it. For the same reason it will also not be possible to compute the exact set of Pareto optimal points $\operatorname*{argMin}\{g(x)\;|\;x\in\mathbb{X}\}$ or the exact set of $\epsilon$-Pareto optimal points $\epsilon\operatorname*{argMin}\{g(x)\;|\;x\in\mathbb{X}\}$ of a convex (but not linear) problem (2), as one could only expect to compute a finitely generated set. Hence, it is hopeless to expect to compute the set $\operatorname{NE}(f,\mathbb{X})$ or $\epsilon\operatorname{NE}(f,\mathbb{X})$ via (5), respectively (6), in Theorems 3.1 and 3.2, exactly. However, below we propose Algorithm 4.7 that computes a finitely generated set $X$ satisfying
@@
as proven in Theorem 4.8. Thus, one obtains an even better approximation of the set of Nash equilibria than the set of all $\epsilon$-approximate Nash equilibria would provide. This will be illustrated in Section 5, see in particular Figure 2 and 4.
@@
  • Assumption 4.2 is satisfied for many games found in the literature (see, e.g., [13, GNEP (21)] and [2, Proposition 3]). Included are also special cases like mixed strategies (without further constraints) as then $\mathbb{X}=[0,1]^{N}$ or, more generally, box constraints.

### v4

@@ 1 Introduction
It is well known that a Nash equilibrium does not need to be Pareto optimal and a Pareto optimal point does not need to be a Nash equilibrium (when considering the corresponding naïve Pareto optimization problem, see e.g. [6] for details). However, in [6], it was shown that the set of Nash equilibria can be equivalently characterized by the set of Pareto optimal points of a specific vector optimization problem with a non-convex ordering cone, or equivalently, by the intersection of the Pareto optimal points of $N$ specific multi-objective problems. This result holds true in general for any possible non-cooperative game without making assumptions on cost functions or constraint sets. In order to use this result for numerical computations, one would need to compute the set of Pareto optimal points for multi-objective problems. So far, algorithms that provide the set of Pareto optimal points exist only for the special case of linear multi-objective optimization, see e.g. [1, 25, 24]. This restricts computational methods of finding the set of Nash equilibria via that characterization to linear games so far.
@@
This result provides some advantages: It is guaranteed that each element of interest, namely each equilibrium point, is contained in the computed set. Furthermore, each element in the found set is guaranteed to be at least almost at equilibrium, i.e. deviating at most by $\epsilon$ from the cost provided by a true equilibrium point. Additionally to the Lipschitz continuity for the jointly convex cost functions, we distinguish between two different assumptions on the structure of the shared constraint set: In the first case we assume that the shared constraint set is a polytope. In the second case we consider for each player $i$ a convex constraint set that is independent from the constraints of the other players. For both cases, the proposed algorithm is proven to work correctly. Computational examples are provided which illustrate the sandwich result.
@@ 2 Definitions
Let us consider the following non-cooperative shared constraint games for $N\geq 2$ players. Each player $i$ (for $i=1,...,N$) considers strategies in the linear space $\mathcal{X}_{i}=\mathbb{R}^{n_{i}}$, where $n_{i}\in\mathbb{N}$, and where we do not impose any condition that $\mathcal{X}_{i}$ and $\mathcal{X}_{j}$ are equal. Furthermore, each player $i$ has a cost function $f_{i}:\prod_{j=1}^{N}\mathcal{X}_{j}\to\mathbb{R}$ she seeks to minimize. This cost $f_{i}(x)$ for $x\in\prod_{j=1}^{N}\mathcal{X}_{j}$ may depend on player $i$’s strategy $x_{i}\in\mathcal{X}_{i}$ as well as the strategy chosen by all other players $x_{-i}\in\prod_{j\neq i}\mathcal{X}_{j}$. The vector of cost functions is denoted by $f=(f_{1},...,f_{N})$ with $f:\prod_{i=1}^{N}\mathcal{X}_{i}\to\mathbb{R}^{N}$. Assume further a shared constraint set $\mathbb{X}\subseteq\prod_{j=1}^{N}\mathcal{X}_{j}$ for the joint strategy $x\in\prod_{j=1}^{N}\mathcal{X}_{j}$ of all players. This shared constraint condition is common in the literature (see, e.g., [21, 4, 12]). In a non-cooperative shared constraint game each player $i$ minimizes her cost function given all other players fix their strategies $x_{-i}^{*}\in\prod_{j\neq i}\mathcal{X}_{j}$. That is, for all $i=1,...,N$ the following optimization problem is considered
@@
Non-cooperative shared constraint games are a special type of generalized games as not only the cost function, but also the constraint set of player $i$ in optimization problem (1) can depend on the strategy $x_{-i}^{*}$ of all other players. In the notation of [12], in which such problems are explicitly called “generalized Nash games with shared constraints”, the constraint for player $i$ can be denoted by $K_{i}(x_{-i}^{*}):=\{x_{i}\in\mathcal{X}_{i}\;|\;(x_{i},x_{-i}^{*})\in\mathbb{X}\}$ so that problem (1) can be written as $x_{i}^{*}\in\operatorname{argmin}\{f_{i}(x_{i},x_{-i}^{*})\;|\;x_{i}\in K_{i}(x_{-i}^{*})\}$. Shared constraints can be used e.g. for system-wide constraints, for details see [12]. Notably, shared constraint games encompass classical games in which the players interact only through the cost functions $f_{i}$. In our notation, a classical game can be encoded by the box-type shared constraint $\mathbb{X}=\prod_{j=1}^{N}\mathbb{X}_{i}$ for $\mathbb{X}_{i}\subseteq\mathcal{X}_{i}$ for every player $i$.
@@
When numerical methods are used, one has to deal with approximation errors. Therefore, we will now introduce the notion of approximate Nash equilibria. In a Nash equilibrium, no player has an incentive to change his behavior. An approximate Nash equilibrium allows the possibility that a player may have a small incentive to deviate. They are also needed for, e.g., sensitivity analysis of Nash equilibria, see [5], or can be interpreted as Nash equilibria to some perturbed preferences [8].
@@
[20, Section 2.6.6] Let $\epsilon>0$. The joint strategy $x^{*}\in\mathbb{X}$ is called an $\bm{\epsilon}$-approximate Nash equilibrium for the game if, for any player $i$ it holds $f_{i}(x_{i},x_{-i}^{*})+\epsilon\geq f_{i}(x^{*})$ for any $x_{i}\in\mathcal{X}_{i}$ such that $(x_{i},x_{-i}^{*})\in\mathbb{X}$. The set of all $\epsilon$-approximate Nash equilibria will be denoted by $\epsilon\operatorname{NE}(f,\mathbb{X})$. That is,
@@
Let us now introduce the notion of Pareto optimality from multi-objective optimization, which also plays a fundamental role in economics as it is related to the interpersonal incomparability of utility. Consider a linear space $\mathcal{X}$ as well as the space $\mathbb{R}^{m}$ with the natural ordering cone +$\mathbb{R}^{m}_{+}$. Recall that a set $C\subseteq\mathbb{R}^{m}$ is called a cone if $\alpha C\subseteq C$ for all $\alpha\geq 0$. The convex cone +$\mathbb{R}^{m}_{+}$ introduces a partial order $\leq$ on $\mathbb{R}^{m}$ via $x\leq y\iff y-x\in\mathbb{R}^{m}_{+},$ for $x,y\in\mathbb{R}^{m}$. A multi-objective optimization problem is a problem of the form
@@
We use the notation $g[\mathbb{X}]:=\{g(x)\;|\;x\in\mathbb{X}\}\subseteq\mathbb{R}^{m}$ for the image of the feasible set $\mathbb{X}$. Note that the set $\operatorname*{argMin}\{g(x)\;|\;x\in\mathbb{X}\}$ can equivalently be characterized as the set of feasible points $x^{*}$ that map to minimal elements of $g[\mathbb{X}]$. That is, $x^{*}\in\mathbb{X}$ is Pareto optimal if, whenever $g(x)\leq g(x^{*})$ for some $x\in\mathbb{X}$, then $g(x^{*})\leq g(x)$ holds, see [9, Def. 3.1(c), Def. 4.1(a) and Def. 7.1(a)].
Denote by $\mathcal{P}:=\operatorname{cl}(g[\mathbb{X}]+\mathbb{R}^{m}_{+})$ the upper image of problem (2). The boundary of the set $\mathcal{P}$ contains the so called efficient frontier, which is the collection of functional values $g(x^{*})$ of all Pareto optimal points $x^{*}$ of problem (2). Numerical methods typically only compute approximations of $\mathcal{P}$, except for the case of linear multi-objective optimization problems, where the set $\mathcal{P}$ as well as the set of all Pareto optimal points can be computed exactly. This can be seen e.g. for the case of convex problems, where the algorithms in [14] result in a polyhedral inner approximation of $\mathcal{P}$ for a given error level $\epsilon>0$ and a fixed direction $c\in\mathbb{R}^{m}_{+}\setminus{\{0\}}$. This motivated the following definition of $\epsilon$-Pareto optimal points, which are feasible points $x^{*}\in\mathbb{X}$ whose image values $g(x^{*})$ are only in $\epsilon$-distance (in direction $c$) to the efficient frontier.
@@
[13][7, Def. 2.2] Let $c\in\mathbb{R}^{m}_{+}\setminus{\{0\}}$. An element $x^{*}\in\mathbb{X}$ is called $\bm{\epsilon}$-Pareto optimal with respect to $c$ for problem (2) if
@@
Definition 2.4 holds for an arbitrary choice of direction $c\in\mathbb{R}^{m}_{+}\setminus{\{0\}}$. However, as we will see in the following section, for a specific vector optimization problem a connection between the set of $\epsilon$-Pareto optimal points and the set $\epsilon\operatorname{NE}(f,\mathbb{X})$ for some $N$-player game can be explicitly stated for the fixed direction $\bar{c}:=(0,...,0,1)^{\top}\in\mathbb{R}^{m}_{+}\setminus{\{0\}}$. Therefore we will, for the remainder of this paper, consider the fixed direction $\bar{c}$ and denote the set of all $\epsilon$-Pareto optimal points with respect to $\bar{c}$ for problem (2) by
@@
[7] $x^{*}\in\mathbb{X}$ is called weakly Pareto optimal for problem (2) if $\big{(}g(x^{*})-\operatorname{int}\mathbb{R}^{m}_{+}\big{)}\cap g[\mathbb{X}]=\emptyset.$ Let $c\in\mathbb{R}^{m}_{+}\setminus{\{0\}}$. An element $x^{*}\in\mathbb{X}$ is called weakly $\epsilon$-Pareto optimal with respect to $c$ for problem (2) if $\big{(}g(x^{*})-\epsilon c-\operatorname{int}\mathbb{R}^{m}_{+}\big{)}\cap g[\mathbb{X}]=\emptyset.$
@@ 3 Equivalence of Nash and Pareto and their approximations
That is, we consider problem (2) with objective function $g(x)=(x_{-i},-x_{-i},f_{i}(x))$. Note that the ordering cone of problem (3) is +$\mathbb{R}^{m_{i}}_{+}$ with $m_{i}=2\sum_{j=1,\\ j\neq i}^{N}n_{j}+1$. This objective space dimension can be further reduced by considering the equivalent convex multi-objective optimization problem
@@
with natural ordering cone in dimension $\sum_{j=1,\\ j\neq i}^{N}n_{j}+2$, where $x_{jk}$ denotes the $k$-th element of the vector $x_{j}$, for $j\in\{1,...,N\}$ and $k\in\{1,...,n_{j}\}$. Hence, $s:=-\sum_{j=1,\\ j\neq i}^{N}\sum_{k=1}^{n_{j}}x_{jk}$ is the sum over all components of the vector $-x_{-i}$. The equivalence holds as minimizing component-wise the vector $(x_{-i},-x_{-i})$ gives the same minimizers (being the set of all $x_{-i}$) as minimizing the vector $(x_{-i},s)$. This dimension reduction is important for computational aspects and should be kept in mind. For purely aesthetical reasons and since it simplifies explanations we will work with problem (3) instead of problem (4) in our exposition. A further dimension reduction by using a non-standard ordering cone is discussed in Remark 4.12.
@@
The above theorem was given in [6] in a more general framework and in a different formulation using a non-convex ordering cone (see [6, Th. 2.6, Cor. 2.8]). We provide here the above intuitive and simplified formulation that was already mentioned in Remark 3.5 in [6] for the linear case, but holds of course also in general. A short remark connecting the algorithm proposed in this paper to the problem formulation using the non-convex ordering cone is given in Remark 4.12.
@@
Theorems 3.1 and 3.2 can also be stated:
  • for generalized games with individual constraint sets $\mathbb{C}_{i}\subseteq\prod_{j=1}^{N}\mathcal{X}_{j}$ for $i\in\{1,...,N\}$, i.e., when problem (1) in Definition 2.1 is replaced by $x_{i}^{*}\in\operatorname*{arg\,min}\{f_{i}(x_{i},x_{-i}^{*})\;|\;(x_{i},x_{-i}^{*})\in\mathbb{C}_{i}\}$. Then, the constraints sets $\mathbb{X}$ in the Pareto problems in Theorems 3.1 and 3.2 have to be replaced by the individual constraint sets $\mathbb{C}_{i}$ for $i\in\{1,...,N\}$.
  • for general linear spaces $\mathcal{X}_{i}$ for the strategies. Then, the ordering cone of the corresponding vector optimization problem has to be adapted accordingly, see Theorem 2.6 in [6].
@@ 4 Convex games
The aim of this paper is to develop an algorithm that uses representations (5) and (6) to approximate the set of Nash equilibria of a non-cooperative game. To do so, we will focus for the remaining part of this paper on convex games satisfying the following assumption.
@@
As mentioned above in issue (i), we try to cover the set of all Pareto optimal points and therefore make additional assumptions on the structure of the constraint set $\mathbb{X}$. In the following, we will consider constraints sets $\mathbb{X}$ that are polytopes, whereas in Remark 4.13, we consider the case where each player $i$ has an independent convex constraint set $\mathbb{X}_{i}\subseteq\mathcal{X}_{i}$. Problem (iii) can be handled by a small modification of the algorithm in [14] that is possible because of the particular structure of the objective function of problem (3) being linear in all but the last component and the directions $\bar{c}_{i}$ being zero in these linear components.
Let us now discuss problem (ii) in detail. The reason why approximations are considered for convex problems is that it is in general not possible to generate the exact upper image by finitely many points or to obtain an analytic expression for it. For the same reasons, only approximations of set of Pareto optimal points $\operatorname*{argMin}\{g(x)\;|\;x\in\mathbb{X}\}$ respectively of the set of $\epsilon$-Pareto optimal points $\epsilon\operatorname*{argMin}\{g(x)\;|\;x\in\mathbb{X}\}$ of a convex (but not linear) problem (2) will be considered. We will focus here on polyhedral approximations as they are finitely generated.
Hence, we cannot expect to compute the sets $\operatorname{NE}(f,\mathbb{X})$ or $\epsilon\operatorname{NE}(f,\mathbb{X})$ via (5), respectively (6), exactly. However, below we propose Algorithm 4.8 that computes a finitely generated set $X$ satisfying
@@
as proven in Theorem 4.9. Thus, one obtains an even better approximation of the set of Nash equilibria than the set of all $\epsilon$-approximate Nash equilibria would provide. This will be illustrated in Section 5, see in particular Figure 1 and 3.
@@
  • Assumption 4.2 is satisfied for many games found in the literature (see, e.g., [16, GNEP (21)] and [2, Proposition 3]). Included are also special cases like mixed strategies (without further constraints) as then $\mathbb{X}=[0,1]^{N}$ or, more generally, box constraints.